### v1

# Optimizing Inventory Placement for a Downstream Online Matching Problem

[…]
Optimizing Inventory Placement for a Downstream Online Matching Problem
[…]
We study the inventory placement problem of splitting $Q$ units of a single item across warehouses, in advance of a downstream online matching problem that represents the dynamic fulfillment decisions of an e-commerce retailer. This is a challenging problem both in theory, because the downstream matching problem itself is computationally hard, and in practice, because the fulfillment team is constantly updating its algorithm and the placement team cannot directly evaluate how a placement decision would perform.
We compare the performance of three placement procedures based on optimizing surrogate functions that have been studied and applied: Offline, Myopic, and Fluid placement. On the theory side, we show that optimizing inventory placement for the Offline surrogate leads to a $(1-(1-1/d)^{d})/2$-approximation for the joint placement and fulfillment problem. We assume $d$ is an upper bound on how many warehouses can serve any demand location and that stochastic arrivals satisfy either temporal or spatial independence. The crux of our theoretical contribution is to use randomized rounding to derive a tight $(1-(1-1/d)^{d})$-approximation for the integer programming problem of optimizing the Offline surrogate. We use statistical learning to show that rounding after optimizing a sample-average Offline surrogate, which is necessary due to the exponentially-sized support, does indeed have vanishing loss.
On the experimental side, we extract real-world sequences of customer orders from publicly-available JD.com data and evaluate different combinations of placement and fulfillment procedures. Optimizing the Offline surrogate performs best overall, even compared to simulation procedures, corroborating our theory.
1 Introduction
The supply chain of an e-commerce retailer is a large and complicated system whose overall operation results from an extensive stream of interdependent decisions. These decisions range all the way from long-term strategic decisions, including network design and sourcing, to mid-term tactical decisions, including inventory replenishment and placement, to real-time operational decisions, including order fulfillment and last-mile delivery (Chen and Graves 2021). Clearly, the optimality of upstream decisions depends on how downstream decisions will be made. However, jointly optimizing the different stages is computationally difficult, and moreover, the sheer size of modern tech companies inevitably leads to a hierarchical structure where each decision is controlled by a separate team. Therefore, a common approach for upstream teams is to abstractly model the downstream dynamics, or make simplifying assumptions on how downstream teams will behave. The ideal way for teams to coordinate, and communicate updates in their algorithm so that other teams can adjust accordingly, is a fascinating problem for both theory and practice.
In this paper we focus on the interaction between the two e-commerce decisions that lie in the most operational end of the spectrum. We study the problem of an e-commerce retailer faced with making inventory placement decisions followed by on-the-fly fulfillment decisions. The e-commerce retailer manages a network of warehouses that can deliver goods to several different last-mile delivery hubs. Before orders start to arrive, the e-tailer must decide how to split an incoming purchase of inventory units across the multiple warehouses it manages. This corresponds to the inventory placement problem, decided by the upstream team. Afterward, a downstream team will face a stream of customer orders during a time horizon, which is random and therefore not known in advance. Whenever an order is made, the downstream team must decide from which warehouse to deliver the product, depleting a unit of inventory and collecting a reward. This sequential decision-making problem corresponds to the fulfillment problem. The goal of the e-commerce retailer is to maximize the expected reward collected during the time horizon by deciding how to place inventory and how to fulfill demand. These two decisions are tightly linked together. On one hand, fulfillment decisions depend on the current inventory levels at each warehouse, which in turn depend on the initial inventory placement. On the other hand, the optimal inventory placement depends on how fulfillment decisions are going to be made. Despite this connection, only recently has there been increased attention on how these decisions interact and how to make them in a joint or coordinated manner (Govindarajan et al. 2021b, Bai et al. 2022a, Chen et al. 2022, Arlotto et al. 2023).
Fulfillment as an isolated problem has been well-studied by the operations research and computer science communities (Xu et al. 2009, Acimovic and Graves 2015, Jasin and Sinha 2015). The main tool of analysis for this problem is modeling it as an online bipartite matching problem with stochastic inputs (Feldman et al. 2009, Alaei et al. 2012, Brubach et al. 2020). These models take the initial inventories at different warehouses as input, and the policies developed aim to obtain a reward as close as possible to the best-achievable reward with that inventory. On the other hand, the inventory placement problem has been less studied. Here, a fixed quantity of an item that has already been purchased is to be distributed among a set of warehouses. A fundamental difference between the placement and fulfillment problems is that there are no downstream decisions after fulfillment is done, whereas the effectiveness of an inventory placement directly depends on the fulfillment policy being deployed. For a fixed inventory placement, finding the optimal fulfillment policy can be computationally challenging (Papadimitriou et al. 2021), which means that even evaluating the potential value of a proposed placement is hard.
To overcome this difficulty, an approach taken by the literature and also in practice is to (exactly or approximately) optimize a surrogate function that approximates the value of the optimal fulfillment policy. The following surrogates and corresponding placement procedures have been considered.
  1. 1.
    Fluid Placement: optimizes inventory placement for a fluid relaxation of the problem, in which the total demand of each type is assumed to be deterministic and equal to its expected value. This procedure only requires knowing the first moments of demand distributions and is also the least computationally expensive. It has been used to prove some approximation guarantees for the joint placement and fulfillment problem (Bai et al. 2022b, Chen et al. 2022).
  2. 2.
    Offline Placement: optimizes the reward obtained by a fulfillment policy that knows the total demand for each type in the given random scenario in advance and thus makes hindsight optimal fulfillment decisions. Govindarajan et al. (2021b), DeValve et al. (2023b) has advocated for this procedure in experiments. This procedure requires the distributional knowledge of total demands and the ability to enumerate over it, which can be difficult. Therefore, often it is approximated by the sampling of demand scenarios.
  3. 3.
    Myopic Placement: makes the simplifying assumption that the downstream fulfillment decisions will be made according to a greedy policy that maximizes immediate rewards, ignoring the opportunity cost of depleting a unit of inventory at a given warehouse. This is an approach present in the lines of work of Acimovic and Graves (2017), Chen (2017), which presents it as a natural first step. Unlike the first two procedures, the surrogate optimized here is rather pessimistic about the sophistication of the downstream fulfillment team. The output of myopic fulfillment depends on the order in which demand arrives, so this information has to be incorporated when optimizing inventory placement for this surrogate, which is usually done through simulation.
Despite the aforementioned work, there does not seem to be clarity on which placement procedure(s) is best. Fluid Placement comes with theoretical guarantees, but is shown to perform substantially worse than Offline Placement in real-world experiments (DeValve et al. 2023b). Offline Placement on the other hand comes with no theoretical guarantees, and the number of samples required for it to prescribe a good decision is also unclear. Finally, Myopic Placement is the most realistic in terms of not expecting the fulfillment team to foretell demand in advance, but the speed and stability of simulation procedures is questionable.
In this paper, we take a deep dive into understanding how these three common approaches for optimizing inventory placement compare. Our contributions come both from deriving theoretical results and from conducting experiments based on publicly-available real-world data. On the theoretical side, we show that when coupled with an appropriate, good-quality fulfillment policy, Offline Placement offers constant-factor guarantees for the joint placement and fulfillment problem. To the best of our knowledge, we are the first work to establish this kind of guarantee using Offline Placement as the inventory placement procedure. Moreover, our guarantee is strictly better than the best guarantees obtained by using Fluid Placement for the inventory placement, complementing the empirical findings of DeValve et al. (2023b). This result formalizes the notion that Offline Placement, a procedure that can be interpreted as being optimistic about the quality of the downstream fulfillment decisions, is desirable when the fulfillment policy being deployed is of high quality. On the experimental side, we extensively and thoroughly compare the performance of different placement procedures on the JD.com dataset from the 2020 MSOM Data-driven Research Challenge (Shen et al. 2020). Our results suggest that Myopic Placement has superior performance to Fluid Placement, supporting the approach of Acimovic and Graves (2017). The comparison between Offline and Myopic Placement is a closer one however; which one outperforms the other is dependent on the fulfillment policy being used and the instance parameters. Our results suggest that Offline Placement has a more robust performance when coupled with a high-quality fulfillment policy, which corroborates our theoretical finding.
1.1 Theoretical Results and Techniques
In Section 4, we study the joint placement and fulfillment problem under two demand models, which we refer to as Temporal Independence and Spatial Independence. The former corresponds to the staple model studied in online matching with stochastic input (Alaei et al. 2012, e.g.), where the time horizon is split into $T$ time steps, and during each time step at most one order will arrive from a specific geographical region (associated with a last-mile delivery hub). As the name suggests, the geographical origin of demand is independent over time, which does not capture settings where arrivals from a location beget more arrivals to follow. In the second model, recently introduced by Aouad and Ma (2022), the total demand from each last-mile hub is sampled independently, and the order in which the demand arrives is decided by an adversary that knows how fulfillment decisions are going to be made and optimizes against it.
We show that under either demand model, when paired with an appropriate fulfillment procedure (depending on the demand model), Offline Placement collects a reward that is at least $(1-(1-1/d)^{d})/2$ times what the optimal joint placement and fulfillment procedures would collect, where $d$ is the maximum number of warehouses that can fulfill a single demand type (Theorem 4.13 in Subsection 4.4). To our knowledge, this is the first result showing that Offline Placement achieves a constant-factor guarantee for the joint placement and fulfillment problem. Here, $d$ is a measure of fulfillment flexibility, which can also be interpreted as an upper bound on the number of warehouses with positive reward for serving any given demand location.
Our guarantee is strictly better than the $(1-1/e)/2$-approximation implied by Bai et al. (2022b) for Fluid Placement (their guarantee is stated as 1/4 and holds in a generalized setting; it improves to $(1-1/e)/2$ in the setting stated here). Our result also provides theoretical justification for why Offline appears to outperform Fluid Placement in the experiments of DeValve et al. (2023b). As $d\to\infty$, our guarantee decreases to approach their guarantee of $(1-1/e)/2$, which is obtained by using submodular optimization to approximately solve Fluid Placement within a factor of $1-1/e$. However, the submodular optimization approach cannot improve (see Section 7) as $d$ gets small, which is often the case in practice—$d=2$ in the celebrated long-chain network (Jordan and Graves 1995) or the JD.com network (Shen et al. 2020). This is why we take a randomized rounding approach instead, and moreover use it to analyze Offline Placement, since we cannot use integrality properties about demand in Fluid Placement. Another benefit of using Offline Fulfillment is that in some arrival models of online matching, the fluid relaxation is too loose to yield constant-factor competitive ratios for online algorithms, as is the case in the Spatial Independence model (Aouad and Ma 2022, Proposition 1).
The crux of our theoretical contribution is to use randomized rounding to provide an approximation algorithm for placement optimization under the Offline surrogate that improves when $d$ is small. This is a fundamental integer programming problem related to matching that is surprisingly non-trivial and, to the best of our knowledge, previously unsolved. The problem is as follows. We are given a bipartite graph with warehouses $i\in[n]:=\{1,\ldots,n\}$, demand types $j\in[m]$, and rewards $r_{ij}\geq 0$ such that |$|\{i:r_{ij}>0\}|\leq d$ for all $j$. A demand vector $D=(D_{1},\ldots,D_{m})$ is drawn from an arbitrary distribution over $\mathbb{Z}_{\geq 0}^{m}$. The problem is to decide an inventory placement $x=(x_{1},\ldots,x_{n})\in\mathbb{Z}_{\geq 0}^{n}$, subject to a constraint $x_{1}+\cdots+x_{n}=Q$ where the total inventory $Q$ has been decided already, to maximize the expected value of
$\mathsf{OFF}(x,D)=\max_{y\geq 0}\quad$ $\sum_{i\in[n]}\sum_{j\in[m]}y_{ij}r_{ij}$
$\mathrm{s.t.}\quad$ $\sum_{i\in[n]}y_{ij}\leq D_{j}\quad\forall j\in[m],$
$\sum_{j\in[m]}y_{ij}\leq x_{i}\quad\forall i\in[n],$
over the randomly-drawn demand vector $D$. Assuming for now that the distribution of $D$ has small support, the fractional relaxation of the placement problem,
$\max_{x\in\mathbb{R}^{m}_{\geq 0},\sum_{i}x_{i}=Q}\mathbb{E}_{D}[\mathsf{OFF}(x,D)],$ (1)
can be formulated as an LP by enumerating over the possible realizations of $D$. It is important to note that the optimal solution of (1) can indeed be fractional even though the bipartite matching LP for any fixed $D$ is integral. In fact, we show that the value of (1) can shrink by a factor of $1-(1-1/d)^{d}$ when $x$ is constrained to lie in $\mathbb{Z}^{n}_{\geq 0}$ (Lemma 4.7 in Subsection 4.2).
Our main result is to show that this factor is in fact tight, i.e. any fractional placement $x$ can be rounded into an integer placement $R(x)=(R_{1}(x),\ldots,R_{n}(x))\in\mathbb{Z}_{\geq 0}^{n}$ such that $\mathbb{E}_{D}[\mathsf{OFF}(R(x),D)]\geq(1-(1-1/d)^{d})\mathbb{E}_{D}[\mathsf{OFF}(x,D)]$ (Theorem 4.3 in Subsection 4.2). We develop a randomized rounding specifically for this problem based on applying the fundamental rounding procedure of Gandhi et al. (2006) in two iterations, which requires new techniques as we describe in Subsection 4.2.
We combine with a 1/2-competitive online algorithm relative to the offline relaxation in either the Temporal or Spatial Independence models to get an approximation ratio of $(1-(1-1/d)^{d})/2$, with this last step applying the three-step approximation framework of Bai et al. (2022b, a). However, our use of the offline relaxation assumed that we could enumerate over the support of $D$, whereas in the Temporal and Spatial Independence arrival models, the support of $D$ has size exponential in the time horizon and number of demand types, respectively. Nonetheless, we show that given $K$ independent and identically distributed (IID) samples, we can still get an approximation ratio of $(1-(1-1/d)^{d})/2-O(Q\sqrt{(n\log n)/K})$. Offline Placement would solve a sample-average approximation of (1) using the $K$ samples, which has size polynomial in $K$, and we use Rademacher complexity and a vector contraction inequality (Maurer 2016) to show that the sample-average approximation provides a uniform approximation of $\mathbb{E}_{D}[\mathsf{OFF}(x,D)]$ over the continuous feasible region $\{x\in\mathbb{R}^{m}_{\geq 0}:\sum_{i}x_{i}=Q\}$ (Theorem 4.8 in Subsection 4.3). Although this argument is simple in hindsight, we remark that a typical analysis of sample-average approximation for optimization problems (Shapiro and Nemirovski 2005, e.g.) does not go through uniform convergence, and would lose an additional factor of 2 in the approximation ratio instead of losing a factor that is vanishing as $K\to\infty$.
1.2 Experimental Results on Public Data
In Section 5, we complement our theoretical results with experiments using the JD.com data available from the 2020 MSOM Data-driven Research Challenge (Shen et al. 2020). These experiments are inspired by the ones conducted in DeValve et al. (2023b). We use this dataset to evaluate how different placement and fulfillment procedures perform when deployed together in this particular JD.com network structure. The main feature we extract from the data, and what makes these experiments different from our theoretical settings, is the exact arrival sequences of demand. Indeed, our theoretical results are worst-case guarantees over all possible demand distributions, whereas in the experiments we take a particular instance from practice with actual demand sequences that were realized. The theoretical settings also impose independence assumptions on the stochastic demand process, which is not the case in the dataset. Moreover, the theoretical results assume that the demand distribution is known and we can generate IID samples. In our experiments, we use past demand data as samples for our placement and fulfillment procedures and evaluate them out-of-sample on future demand, where there may have been distributional shifts in reality.
We evaluate four different placement procedures: Offline Placement, Myopic Placement, Fluid Placement, and Proportional Placement, a simple procedure that places inventory proportional to expected demand. The fulfillment procedures evaluated are Myopic Fulfillment and several variants of linear programming-based policies that use the shadow prices of their constraints to approximate and incorporate the opportunity costs of depleting inventory from specific warehouses. The variants depend on the linear program being solved (fluid or offline) and whether they include re-solving or not.
Consistent with DeValve et al. (2023b), we find that the fulfillment policy which uses the shadow prices of the offline linear program and re-solves periodically has the best performance. Moreover, when this fulfillment policy is deployed, Offline Placement achieves the best performance among the four placement procedures we benchmark. This result is robust to varying load factors (ratio between expected demand and number of inventory units) and edge rewards. This again backs the notion that deploying Offline Placement leads to a good performance given that the fulfillment policy is of high quality. We also find that Myopic Placement can be a desirable placement procedure to deploy if edge rewards are similar among each other, and thus the main cause for losing reward is lost sales; or of course, if Myopic Fulfillment will indeed be the fulfillment policy. Overall, our experiments suggest that using a placement procedure that takes into account the uncertainty of the demand (Offline Placement or Myopic Placement) is consistently better than using deterministic approximations that use average demand as input (Fluid Placement or Proportional Placement)—and this is despite the fact that the more sophisticated Offline and Myopic Placement procedures are more likely to overfit to the training data.
Table 1 presents a summary of how the three placement procedures of interest compare and how our results contribute to this understanding.
Placement Procedure Offline Placement $(\mathsf{OFF})$ Myopic Placement $(\mathsf{MYO})$ Fluid Placement $(\mathsf{FLU})$
Information Required
Medium (all moments of aggregate demand for each type)
High (exact demand arrival sequence)
Low (first moment of aggregate demand for each type)
Computation Time
Requires sampling
Requires simulation
Fastest (still no poly-time algorithm)
Surrogate Approximation
$1-(1-1/d)^{d}$ $\boldsymbol{(*)}$
(LP and rounding)
$1-1/e$ (Bai et al. 2022b)
(Submodular optimization)
Overall Approximation
$(1-(1-1/d)^{d})/2$ $\boldsymbol{(*)}$
(Under both demand models)
$(1-1/e)/2$ (Bai et al. 2022b)
(Only under Temporal Independence)
Empirical Justification
Beats $\mathsf{FLU}$ (DeValve et al. 2023b)
Beats $\mathsf{MYO}$ given a good fulfillment policy $\boldsymbol{(*)}$
Beats $\mathsf{FLU}$ $\boldsymbol{(*)}$
Table 1: Comparison between three placement procedures. Entries accompanied by a $\boldsymbol{(*)}$ represent contributions of this paper.
2 Further Related Work
Joint inventory placement and fulfillment. Chen et al. (2022) studies the joint inventory placement and fulfillment problem with multiple items, obtaining a $(1-1/e)/4$ worst-case guarantee, and demonstrating its applicability at Anheuser Busch InBev (ABI). Bai et al. (2022b) study the joint inventory placement, promise, and fulfillment problem with multiple items and warehouse capacity constraints, obtaining a $1/4$ worst-case guarantee. Their proposed inventory placement uses submodular optimization techniques to obtain an approximate solution for a fluid linear program. Bai et al. (2022a) studies the more general problem where the downstream decisions correspond to deciding an assortment, obtaining a $(1-1/e)/4$-approximation. Govindarajan et al. (2021b) study the joint distribution and placement problem in the context of an omni-channel retailer. They provide an inventory placement heuristic that is, as the one we analyze, based on solving the optimal inventory placement problem assuming fulfillment will be hindsight-optimal, although they do not provide theoretical guarantees. Arlotto et al. (2023) study the single-item joint inventory placement and fulfillment problem through the lens of regret minimization. They obtain sublinear regret for two state-of-the-art fulfillment policies, assuming that the initial inventory is optimal for the specific policy. All of these models use the classic Temporal Independence demand model, whereas the results we obtain also hold in the more recently introduced Spatial Independence demand model.
Other coordinated decisions in e-commerce retail. The recent trend of studying coordinated decisions in e-commerce retail goes beyond inventory placement and fulfillment. Lei et al. (2018) study the problem of jointly pricing and fulfillment for different items of a catalog during a fixed time horizon. Lei et al. (2022) study the more general problem of joint product framing and order fulfillment under the multinomial logit model. They solve a deterministic approximation for the stochastic control problem and develop a randomized rounding scheme for this solution. Jasin et al. (2022) studies a joint inventory replenishment, pricing, and fulfillment problem with multiple stores and warehouses. They develop a Lagrangian based heuristic that achieves sub-linear regret in the time horizon. DeValve et al. (2023b) primarily studies the effects of flexibility in e-commerce retail supply chains, but their work takes a look at how several aspects interact. Through numerical experiments using real-world data, they study the interaction of fulfillment, placement, and network design. In particular, they find that if the deployed fulfillment policy is myopic, adding flexibility to the network can increase total costs. DeValve et al. (2023a) study the two-stage problem of deciding a network structure followed by fulfillment decisions, which they assume can be made in an offline fashion, and derive constant factor guarantees. Zhao et al. (2023) study an inventory replenishment and fulfillment problem in a network structure that resembles the JD.com network that we study. DeValve (2023) studies, among other problems, the multi-item joint inventory buying, placement, and fulfillment problem with the presence of fixed costs, where orders can contain several items. They show that it is NP-hard to approximate this problem up to a constant factor. They also derive approximation guarantees for a broad range of problems.
Newsvendor Networks. Van Mieghem (1998), Mieghem and Rudi (2002) study a multi-dimensional newsvendor model, closely related to optimizing inventory assuming that fulfillment will be hindsight optimal. The difference with inventory placement is that in the latter, the total number of units, $Q$, has already been decided exogenously. Closely related is Govindarajan et al. (2021a), where they study a multi-location newsvendor network model where the only information known about the joint demand distribution is the mean vector and covariance matrix. They take a distributionally robust approach to find the inventory placement that minimizes the worst-case expected cost of fulfilling demand. For recent developments and an in-depth literature review related to Newsvendor networks, we refer the reader to DeValve and Myles (2023).
3 Theoretical Problem Definition
We study the two-stage problem faced by an e-commerce retailer (e-tailer) who has to make an inventory placement decision at the beginning of a time horizon (first stage), followed by sequential fulfillment decisions during that time horizon (second stage). The e-tailer manages a supply chain network consisting of distribution centers (DCs) $i\in[n]$ that can hold inventory and serve different demand types $j\in[m]$. Demand type $j$ can be thought of as demand originating from a specific district or ZIP code that we call $j$. When type $j$ demand arrives, the e-tailer must decide from which distribution center (if any) that demand is going to be satisfied. If the e-tailer decides to deliver the item from DC $i$, they collect a reward of $r_{ij}\in[0,1]$. (Here, we are assuming that the rewards are bounded and normalized to lie between 0 and 1.) We say that DC $i$ can serve demand of type $j$ if $r_{ij}>0$. For each demand type $j\in[m]$, we define its degree $d_{j}:=|\{i\in[n]:r_{ij}>0\}|$ as the number of DCs that can serve the demand of this type. We further define $d:=\max_{j\in[m]}d_{j}$. Parameter $d$ is an indicator of fulfillment flexibility in the network, and our guarantees will depend on it. In this problem, we focus on a single-item setting.
In the inventory placement (first stage) problem, the e-tailer must distribute an incoming shipment of $Q\in\mathbb{N}$ inventory units across the $n$ distribution centers they manage, where $\mathbb{N}$ are the natural numbers. Formally, the e-tailer must decide on an inventory placement in $\mathcal{X}:=\{x\in\mathbb{N}^{n}:\sum_{i\in[n]}x_{i}=Q\}$. We note that $Q$ is not a decision variable, since we assume that the total quantity to stock has already been determined by an upstream inventory team.
The fulfillment (second stage) problem takes this inventory placement as an input and faces a random stream of demand types. We say that the total number of demand arrivals is $T$, which we call the time horizon and is generally random. We characterize the stream of demand arrivals by a random vector $J=(j_{t})_{t=1}^{T}$, where $j_{t}=j$ if and only if the $t$-th demand request is of type $j$. This stream is unknown to the e-tailer, who only gets to observe the sequence of demand requests as they arrive. We define $D_{j}=\sum_{t=1}^{T}\mathbbm{1}\{j_{t}=j\}$ to be the total number of type $j$ demand that arrives during the time horizon, and we further define $D=(D_{j})_{j\in[m]}$ to be the vector containing these total demands. The e-tailer has an initial inventory $x\in\mathcal{X}$ decided in the first stage, and starts observing the demand sequence one by one. When the $t$-th demand request arrives, the e-tailer observes its origin $j_{t}$ and must irrevocably decide at most one DC $i\in[n]$ with remaining inventory from which to fulfill the request. Once the DC is decided, the e-tailer depletes one unit of inventory, collects a reward $r_{ij}$, and proceeds to observe the next arrival. We say that these online decisions are prescribed by a fulfillment policy $\pi$, and let $\Pi$ denote the class of all admissible fulfillment policies that make decisions without knowing the future.
The overall goal of the e-tailer is to deploy a placement $x$ in conjunction with a fulfillment policy $\pi$ to maximize the expected total reward collected during the time horizon. Our results hold under two different models for the demand arrivals—Temporal and Spatial Independence, detailed below.
Temporal Independence model. In this model, the time horizon $T$ is deterministic and known to the e-tailer. In each time step we will have $j_{t}=j$ with probability $p_{tj}$, where $j_{t}$ is independent of $j_{t^{\prime}}$ if $t\neq t^{\prime}$. These probabilities are also known by the e-tailer. We allow the occurrence that no demand arrives in a given time step. We represent that by a dummy demand type which has zero reward and hence is always rejected. Then, for all $t\in[T]$ we have $\sum_{j\in[m]}p_{tj}=1$. In this context, a policy $\pi$ maps the current time step $t$, the origin of the demand $j_{t}$, and the current inventories to at most one DC for fulfillment. The policy $\pi$ can be chosen based on the instance parameters. We use $\pi(x,J)$ to denote the expected reward collected by a policy $\pi$ starting with inventory $x$ when faced with a demand sequence $J$. We use $\mathsf{ONL}^{\pi}(x)$ to denote the expected reward collected by a policy $\pi$ with starting inventory $x$, where the expectation is taken over the random demand sequence:
$\mathsf{ONL}^{\pi}(x)=\mathbb{E}_{J}[\pi(x,J)].$
We use $\mathsf{OPT}(x)$ to denote the maximum reward that can be obtained by an online policy:
$\mathsf{OPT}(x)=\sup_{\pi\in\Pi}\mathsf{ONL}^{\pi}(x).$
Spatial Independence model. In this model, the total demand $D_{j}$ of each type $j\in[m]$ is sampled from independent distributions $(G_{j})_{j\in[m]}$. Let $G$ be the product distribution $G_{1}\times\cdots\times G_{m}$. Once the total demands are sampled and the e-tailer has chosen its policy, an adversary is allowed to choose the order in which the requests will arrive such that the reward collected by the e-tailer is minimized. Note that the time horizon $T=\sum_{j=1}^{m}D_{j}$ is now random and unknown to the e-tailer. In this model, a fulfillment policy $\pi$ additionally uses the amount of each demand type observed so far to make its decisions, because this can affect the posterior for the future demand of each type. The policy $\pi$ can again be chosen based on the instance parameters. We reuse $\pi(x,J)$ to denote the expected reward collected by a policy $\pi$ starting with inventory $x$ when faced with a demand sequence $J$. Let $\mathcal{J}(D):=\{(j_{t})_{t=1}^{T}\in[m]^{T}:\sum_{t=1}^{T}\mathbbm{1}\{j_{t}=j\}=D_{j}\quad\forall j\in[m]\}$ be the set of arrival orders that the adversary can choose. We again use $\mathsf{ONL}^{\pi}(x)$ to denote the expected reward collected by a policy $\pi$ with starting inventory $x$, but now the expectation is taken over total demands, and the adversary chooses the random order knowing the demand and the policy:
$\mathsf{ONL}^{\pi}(x)=\mathbb{E}_{D\sim G}\left[\min_{J\in\mathcal{J}(D)}\pi(x,J)\right].$
As in the Temporal Independence model, we use $\mathsf{OPT}(x)$ to denote the reward obtained by the best possible policy for a specific instance:
$\mathsf{OPT}(x)=\sup_{\pi\in\Pi}\mathsf{ONL}^{\pi}(x).$
Overall, an instance of our joint placement and fulfillment problem includes the total inventory to be placed $Q$, the sets of DCs and demand types $[n]$ and $[m]$, and the rewards $(r_{ij})_{i,j\in[n]\times[m]}$. In the Temporal Independence model, an instance further includes the demand parameters $T$ and $(p_{tj})_{t,j\in[T]\times[m]}$. On the other hand, in the Spatial Independence model, an instance further includes the demand distribution $G=G_{1}\times\cdots\times G_{m}$. A placement procedure $P$ takes in an instance and outputs an inventory placement $x^{P}\in\mathcal{X}$. Similarly, a fulfillment procedure $F$ takes in an instance along with a placement and outputs a fulfillment policy $\pi^{F}\in\Pi$. For each demand model, our goal is to derive a placement procedure $P$ along with a fulfillment procedure $F$ such that
$\mathsf{ONL}^{\pi^{F}}(x^{P})\geq\gamma\max_{x\in\mathcal{X}}\sup_{\pi\in\Pi}\mathsf{ONL}^{\pi}(x)-\varepsilon=\gamma\max_{x\in\mathcal{X}}\mathsf{OPT}(x)-\varepsilon$
for all instances, with approximation ratio $\gamma\in[0,1]$ being as large as possible, and sampling error $\varepsilon\geq 0$ (to be specified) being as small as possible.
4 Theoretical Results
In this section, we develop our approximation results for the joint inventory placement and fulfillment problem. Our proof is inspired by the three-step approximation framework by Bai et al. (2022b, a). This framework consists of 1) coming up with a surrogate function $f$ such that $f(x)\geq\mathsf{OPT}(x)$ for all $x$, 2) obtaining an inventory placement $x^{\prime}$ such that $f(x^{\prime})\geq\alpha\max_{x\in\mathcal{X}}f(x)$, and 3) coming up with a fulfillment policy such that $\mathsf{ONL}^{\pi}(x^{\prime})\geq\beta f(x^{\prime})$. This way, they can obtain an $\alpha\beta$ approximation by using $x^{\prime}$ and $\pi$:
$\mathsf{ONL}^{\pi}(x^{\prime})\geq\beta f(x^{\prime})\geq\beta\alpha\max_{x\in\mathcal{X}}f(x)\geq\beta\alpha\max_{x\in\mathcal{X}}\mathsf{OPT}(x).$ (2)
In our work, the surrogate function we will use for step 1) is the expected reward obtained by a hindisght optimal or offline fulfillment, which we denote by $\mathsf{OFF}(x)$. For step 2), approximating the problem $\max_{x\in\mathcal{X}}\mathsf{OFF}(x)$, we solve a linear relaxation of the problem and randomly round the solution to obtain an integer placement that achieves a $(1-(1-1/d)^{d})$-approximation of the linear relaxation. Given the nature of the demand models, the size of this surrogate grows exponentially with $T$ (Temporal Independence) or $m$ (Spatial Independence), so it is not clear if we can optimize or even compute the value of this surrogate. To overcome this we resort to sample-average approximation, where we sample a polynomial number of demand realizations and optimize the sample average instead. This extra sampling does not allow us to use the clean chain of inequalities in Equation 2, so here we deviate from the three-step approximation framework and apply some learning theory. Finally, for step 3) we borrow the fulfillment procedures from Alaei et al. (2012), Aouad and Ma (2022) that offer guarantees with respect to offline fulfillment, as formalized below.
Proposition 4.1 (Alaei et al. (2012), Aouad and Ma (2022))
Under either demand model, for any instance and any (integer) placement $x\in\mathcal{X}$, a fulfillment policy $\pi$ can be computed with
$\mathsf{ONL}^{\pi}(x)\geq\frac{1}{2}\mathsf{OFF}(x).$
The rest of this section is organized as follows. In Subsection 4.1 we introduce the expected reward of offline fulfillment as our surrogate function. In Subsection 4.2 we introduce and show the guarantees of our randomized rounding that takes a fractional solution and returns an integer inventory placement. In Subsection 4.3 we use tools from statistical learning theory to show that we can obtain an arbitrarily good solution from our sample-average approximation using a polynomial number of samples. In Subsection 4.4 put all the pieces together and establish our main result.
4.1 Proposed Surrogate
The first step for obtaining our approximation is to propose a surrogate function that upper-bounds $\mathsf{OPT}(x)$ for any $x\in\mathcal{X}$. The surrogate function that we will be using corresponds to the expected value of a hindsight optimal or offline fulfillment. Offline fulfillment does not depend on the order in which demand requests arrive, but solely on the vector $D$ of total demand arrivals per type $j\in[m]$. For a given demand vector $D$, we let $\mathsf{OFF}(x,D)$ denote the reward collected through offline fulfillment with starting inventory $x$. This value corresponds to the optimal value of the following bipartite matching linear program:
$\mathsf{OFF}(x,D)=\max_{y\geq 0}\quad$ $\sum_{i\in[n]}\sum_{j\in[m]}y_{ij}r_{ij}$
$\mathrm{s.t.}\quad$ $\sum_{i\in[n]}y_{ij}\leq D_{j}\quad\forall j\in[m],$ (3)
$\sum_{j\in[m]}y_{ij}\leq x_{i}\quad\forall i\in[n],$ (4)
and the ‘fulfillment decisions’ correspond to the optimal LP variables $y_{ij}$, interpreted as the number of units delivered from DC $i$ to demand type $j$. We then have $\mathsf{OFF}(x)=\mathbb{E}_{D}[\mathsf{OFF}(x,D)]$, where $\mathsf{OFF}(x)$ is averaging over the demand vector $D$. The following result is well-known.
Proposition 4.2
For any demand model, any policy $\pi\in\Pi$ and any starting inventory $x\in\mathcal{X}$, it holds that $\mathsf{OFF}(x)\geq\mathsf{ONL}^{\pi}(x)$.
Before proceeding, we note that both $\mathsf{OFF}(x,D)$ and $\mathsf{OFF}(x)$ are well-defined for a fractional placement $x$ lying in the convex hull of $\mathcal{X}$, denoted by $\mathrm{CH}(\mathcal{X})=\{x\in\mathbb{R}^{n}_{+}:\sum_{i\in[n]}x_{i}=Q\}$, where $\mathbb{R}_{+}$ is the set of all non-negative real numbers. This observation is important since we will be working with the linear relaxation of $\max_{x\in\mathcal{X}}\mathsf{OFF}(x)$.
4.2 Approximating the Surrogate: Randomized Rounding
One possible approach for approximating $\max_{x_{\in}\mathcal{X}}\mathsf{OFF}(x)$ is to use submodular optimization techniques. However, this approach fails to get an approximation that improves for small $d$. Bai et al. (2022b) show that $\mathsf{OFF}(x,D)$ is a submodular function over $x\in\mathcal{X}$ for any fixed $D$. Since submodularity is preserved under a convex combination, $\mathsf{OFF}(x)$ is also a submodular function over $x\in\mathcal{X}$. It is straightforward to verify that $\mathsf{OFF}(x)$ is also a non-decreasing function of $x$, implying that we could obtain a $(1-1/e)$-approximation of $\max_{x\in\mathcal{X}}\mathsf{OFF}(x)$ by constructing an inventory placement greedily (Calinescu et al. 2007). More precisely, the submodular approach would yield an approximation of $(1-(1-1/Q)^{Q})$, where $Q$ is the number of inventory units to be placed. In applications $Q$ is large, so this approximation would be close to $1-1/e$. This factor cannot be further improved even if $d$ is small, as shown in Section 7.
Therefore, our approach for approximating $\max_{x_{\in}\mathcal{X}}\mathsf{OFF}(x)$ is to instead use randomized rounding, which will get an approximation guarantee that improves for small $d$. In particular, we solve the relaxation $\max_{x\in\mathrm{CH}(\mathcal{X})}\mathsf{OFF}(x)$ using linear programming and then establishing a randomized rounding for this problem. Our randomized rounding applies the (star graph case of the) rounding procedure of Gandhi et al. (2006) in two iterations. This procedure receives as input a finite set $A$ and a weight $w_{i}\in[0,1]$ for every $i\in A$. It outputs a vector of random variables $W_{i}\in\{0,1\}$ for all $i$ that satisfies three properties:
  • (P1)
    Marginal Distribution: $\mathbb{E}[W_{i}]=w_{i}$ for all $i\in A$.
  • (P2)
    Degree Preservation: Let $\delta=\sum_{i\in A}w_{i}$. Then $\sum_{i\in A}W_{i}\in\{\lfloor\delta\rfloor,\lceil\delta\rceil\}$ with probability 1.
  • (P3)
    Negative Correlation: For any subset $S\subseteq A$ and for any $b\in\{0,1\}$:
    $\mathbb{P}\left(\bigcap_{i\in S}W_{i}=b\right)\leq\prod_{i\in S}\mathbb{P}(W_{i}=b).$
This randomized rounding will be used to round the fractional components of our placement $x\in\mathrm{CH}(\mathcal{X})$, and its properties will be used in our proof. We feed in the set of DCs, i.e. $A=[n]$, with weights $w_{i}=x_{i}-\lfloor x_{i}\rfloor$ for all $i$. For a fractional solution $x\in\mathrm{CH}(\mathcal{X})$, we will use $R(x)$ to denote the rounded output, with $R_{i}(x)$ being the rounded inventory for DC $i$. We only round the fractional part of the solution and leave the whole part untouched. This means that we will have $R_{i}(x)=\lfloor x_{i}\rfloor+1$ with probability $x^{f}_{i}:=x_{i}-\lfloor x_{i}\rfloor$ and $R_{i}(x)=\lfloor x_{i}\rfloor$ with probability $1-x^{f}_{i}$. In our application, the three properties translate to:
  1. (P1)
    Marginal Distribution: $\mathbb{E}[R_{i}(x)]=x_{i}$ for all $i\in[n]$.
  2. (P2)
    Degree Preservation: $\sum_{i\in[n]}R_{i}(x)=Q$ with probability 1.
  3. (P3)
    Negative Correlation: For any subset $S\in[n]$:
    $\mathbb{P}\left(\bigcap_{i\in S}\{R_{i}(x)=\lfloor x_{i}\rfloor+1\}\right)\leq\prod_{i\in S}\mathbb{P}\left(R_{i}(x)=\lfloor x_{i}\rfloor+1\right)$
    $\text{and}\quad\mathbb{P}\left(\bigcap_{i\in S}\{R_{i}(x)=\lfloor x_{i}\rfloor\}\right)\leq\prod_{i\in S}\mathbb{P}\left(R_{i}(x)=\lfloor x_{i}\rfloor\right).$
We now establish the following guarantee for this rounding procedure on our problem.
Theorem 4.3
For any instance such that the maximum degree across demand types is $d$ and for any fractional inventory placement $x\in CH(\mathcal{X})$, it holds that
$\mathbb{E}_{R}[\mathsf{OFF}(R(x))]\geq\left(1-\left(1-\frac{1}{d}\right)^{d}\right)\mathsf{OFF}(x).$ (5)
Note that $\mathbb{E}_{R}[\mathsf{OFF}(R(x))]=\mathbb{E}_{D,R}[\mathsf{OFF}(R(x),D)]$ and $\mathsf{OFF}(x)=\mathbb{E}_{D}[\mathsf{OFF}(x,D)]$. Therefore, to prove (5), it suffices to show that $\mathbb{E}_{R}[\mathsf{OFF}(R(x),D)]\geq(1-\left(1-\frac{1}{d})^{d}\right)\mathsf{OFF}(x,D)$ for any fixed realization of $D$. To show this for a fixed realization of $D$, we will split each type $j$ into $D_{j}$ identical types with one unit of demand each. After doing so, we end up with the same problem except that now $D_{j}=1$ for all $j\in[T]$ under the new indexing. (Note that we cannot split inventory units at DCs in the same way, because that would change the parameter $d$ in our guarantee.)
[…]
We need to prove several properties before establishing Theorem 4.3. First, we need to show that for every $i\in[n]$, we can find probabilities $(y_{ij}^{{}\mathrm{H}})_{j\in[T]}$ and $(y_{ij}^{{}\mathrm{L}})_{j\in[T]}$ that satisfy the following constraints:
$y_{ij}^{{}\mathrm{H}}x_{i}^{f}+y_{ij}^{{}\mathrm{L}}(1-x_{i}^{f})$ $=y_{ij}$ $\quad\forall j\in[T],$ (6)
$0\leq y_{ij}^{{}\mathrm{L}}\leq y_{ij}^{{}\mathrm{H}}$ $\leq 1$ $\quad\forall j\in[T],$ (7)
$\sum_{j\in[T]}y_{ij}^{{}\mathrm{L}}$ $\leq\lfloor x_{i}\rfloor,$ (8)
$\sum_{j\in[T]}y_{ij}^{{}\mathrm{H}}$ $\leq\lfloor x_{i}\rfloor+1.$ (9)
Equation 6 will make $\mathbb{E}[Y_{ij}]=y_{ij}$ for all $j\in[T]$. Equation 7 is a logical monotonicity and will be important for a correlation argument later on. Equations 8 and 9 will make it so that we can indeed always apply the procedure of Gandhi et al. (2006), regardless of whether $R_{i}(x)=\lfloor x_{i}\rfloor$ or $R_{i}(x)=\lfloor x_{i}\rfloor+1$. We note that naive ways for setting these probabilities such as scaling, sampling, or splitting fail:
  1. 1.
    Scaling: Setting $y^{\mathrm{H}}_{ij}=y_{ij}(\lceil x_{i}\rceil/x_{i}),y^{\mathrm{L}}_{ij}=y_{ij}(\lfloor x_{i}\rfloor/x_{i})$ for each $j$ satisfies all the conditions, but could create values of $y^{\mathrm{H}}_{ij}$ that are greater than 1 and hence invalid probabilities.
  2. 2.
    Proportional Sampling: Another plausible strategy is to sample $R_{i}(x)$ times without replacement from $[T]$, with probabilities proportional to $y_{i1},\ldots,y_{iT}$. However, this does not preserve marginals—take for example $y_{i1}=0.9,y_{i2}=0.6,x_{i}=1.5$. This procedure would induce $y^{\mathrm{H}}_{i1}=1$ and $y^{\mathrm{L}}_{i2}=3/5$, but does not satisfy $0.5y^{\mathrm{H}}_{i1}+0.5y^{\mathrm{L}}_{i2}=y_{i1}$.
  3. 3.
    Further Splitting: Naive methods would work if we could split warehouses so that $x_{i}\leq 1$ for all $i$; however, this would change the value of $d$ that we are parametrizing by.
Given the failure of these naive methods, we instead show fundamentally that a feasible adjustment $(y^{\mathrm{L}}_{ij},y^{\mathrm{H}}_{ij})_{j\in[m]}$ exists, which essentially boils down to applying the simple fact that $\lceil x_{i}\rceil-x_{i}\leq\sum_{j}(\lceil y_{ij}\rceil-y_{ij})$ when $\sum_{j}y_{ij}=x_{i}$.
Lemma 4.4
For all $i\in[n]$, there exist vectors of probabilities $(y_{ij}^{{}\mathrm{H}})_{j\in[T]}$ and $(y_{ij}^{{}\mathrm{L}})_{j\in[T]}$ that satisfy Equations 6, 7, 8 and 9, assuming that $\sum_{j\in[T]}y_{ij}\leq x_{i}$ and $y_{ij}\in[0,1]$ for each $j\in[T]$.
The proof of this lemma is deferred to Subsection 6.1. With the properties from Lemma 4.4, we can establish the following negative correlation for indicators $Y_{ij}$ across $i\in[n]$.
Lemma 4.5
For any $j\in[T]$ and for any subset $S\subseteq[n]$, it holds that
$\mathbb{P}\left(\bigcap_{i\in S}Y_{ij}=0\right)\leq\prod_{i\in S}\mathbb{P}(Y_{ij}=0).$
The proof of this lemma is deferred to Subsection 6.2. Equipped with Lemma 4.5, we now lower-bound the reward that can be obtained from each demand type, in expectation over the randomized placement and assignment.
Lemma 4.6
Fix a realization of the demand vector $D$. It holds for all demand types $j\in[T]$ that
$\mathbb{E}\left[\sum_{i\in[n]}r_{ij}Z_{ij}\right]\geq\left(1-\left(1-\frac{1}{d_{j}}\right)^{d_{j}}\right)\sum_{i\in[n]}r_{ij}y_{ij},$
where the expectation is taken over the randomized rounding $R(x)$ and random assignments determined by $Y_{ij}$ and $Z_{ij}$.
The proof of this lemma is deferred to Subsection 6.3. With these results, we now complete the proof of Theorem 4.3, presented in Subsection 6.4.
We close this subsection by noting that the approximation from Theorem 4.3 is tight.
Lemma 4.7
There exists a family of instances for which
$\frac{\max_{x\in\mathcal{X}}\mathsf{OFF}(x)}{\max_{x\in\mathrm{CH}(\mathcal{X})}\mathsf{OFF}(x)}\leq 1-\left(1-\frac{1}{d}\right)^{d}.$
The proof of this lemma is deferred to Subsection 6.5.
4.3 Approximating the Surrogate: Sample-average Approximation
The rounding results in Subsection 4.2 were contingent on being able to solve the problem $\max_{x\in\mathrm{CH}(\mathcal{X})}\mathsf{OFF}(x)$. However, since $\mathsf{OFF}(x)$ takes an expectation over $D$ and the number of possible realizations of $D$ grows exponentially under either the Temporal or Spatial Independence models, the problem cannot be solved exactly. Instead, we solve a sample-average approximation of the problem $\max_{x\in\mathrm{CH}(\mathcal{X})}\mathsf{OFF}(x)$, and prove that the sample-average approximation is not too sensitive to the samples drawn.
To elaborate, we independently sample $K$ vectors $D^{1},\dots,D^{K}$ for the total demand according to the given demand model. We define
$\widehat{\mathsf{OFF}}(x)=\frac{1}{K}\sum_{k=1}^{K}\mathsf{OFF}(x,D^{k}),$
and solve for $\hat{x}$, an optimal solution to the problem $\max_{x\in\mathrm{CH}(\mathcal{X})}\widehat{\mathsf{OFF}}(x)$. In general, we will use the symbol $\wedge$ to refer to anything that depends on the samples $D^{1},\dots,D^{K}$, and we will use $\mathbb{E}_{\wedge}$ when we are taking an expectation over the random samples. It will also be understood that the sample consists of exactly $K$ demand realizations.
We can solve $\max_{x\in\mathrm{CH}(\mathcal{X})}\widehat{\mathsf{OFF}}(x)$ by solving the following linear program:
$\max_{x\in\mathbb{R}^{n}_{+},y\in\mathbb{R}^{[n]\times[m]\times[K]}_{+}}\quad$ $\frac{1}{K}\sum_{k=1}^{K}\sum_{i\in[n]}\sum_{j\in[m]}r_{ij}y_{ij}^{k}$
$\mathrm{s.t.}\quad$ $\sum_{i=1}^{n}y_{ij}^{k}\leq D_{j}^{k}\quad$ $\forall j\in[m],k\in[K],$
$\sum_{i=j}^{m}y_{ij}^{k}\leq x_{i}\quad$ $\forall i\in[n],k\in[K],$
$\sum_{i=1}^{n}x_{i}=Q.$
Notice that this problem has a polynomial number of variables and constraints if $K$ is polynomial in the instance parameters. By contrast, the problem $\max_{x\in\mathrm{CH}(\mathcal{X})}\mathsf{OFF}(x)$ would have required an exponentially-sized $K$ to capture the support of $D$.
The main result of this subsection, which will allow us to use the sample-average solution $\hat{x}$ to obtain our approximation result, is the following.
Theorem 4.8
For a random sample of $K$ IID demand realizations, we have
$\mathbb{E}_{\wedge}\left[\sup_{x\in\mathrm{CH}(\mathcal{X})}\widehat{\mathsf{OFF}}(x)-\mathsf{OFF}(x)\right]=O\left(Q\sqrt{\frac{n\log n}{K}}\right).$
Theorem 4.8 provides a uniform convergence guarantee on the generalization error of how much $\widehat{\mathsf{OFF}}$ can overestimate the value of any solution $x$. We note that the space of fractional solutions $\mathrm{CH}(\mathcal{X})$ is continuous and infinite; however we can apply a vector contraction inequality after showing the function $\mathsf{OFF}(x)$ to be Lipschitz in $x$. We now state the tools from statistical learning theory required for our result.
We first define the Rademacher complexity of our sample as:
^$\widehat{\mathcal{R}}:=\mathbb{E}_{\sigma}\left[\sup_{x\in\mathrm{CH}(\mathcal{X})}\frac{1}{K}\sum_{k=1}^{K}\sigma_{k}\left(1-\frac{\mathsf{OFF}(x,D^{k})}{Q}\right)\right],$ (10)
where $(\sigma_{k})_{k\in[K]}$ are independent Rademacher random variables, defined as random variables that are either $+1$ or $-1$, with probability $1/2$ each. We note that in (10) we are normalizing all values $\mathsf{OFF}(x,D^{k})$ to lie in [0,1] (all rewards $r_{ij}$ lie in [0,1] and we have only $Q$ units, so the total reward is upper-bounded by $Q$) and subtracting it from 1 to get a loss function, to obtain the following result, whose proof is deferred to Subsection 6.6.
Lemma 4.9
$\mathbb{E}\left[\sup_{x\in\mathrm{CH}(\mathcal{X})}\widehat{\mathsf{OFF}}(x)-\mathsf{OFF}(x)\right]\leq 2Q\mathbb{E}_{\wedge}\left[\widehat{\mathcal{R}}\right].$
The following two results that we directly quote will help us upper-bound the Rademacher complexity, thus giving an upper bound to the right-hand side in Lemma 4.9.
Proposition 4.10 (Corollary 4 from Maurer (2016))
For functions $(h_{k}(x))_{k=1}^{K}$ that are $L$-Lipschitz on $x$ in the $\ell_{2}$ norm, we have
$\mathbb{E}_{\sigma}\left[\sup_{x\in\mathrm{CH}(\mathcal{X})}\sum_{k=1}^{K}\sigma_{k}h_{k}(x)\right]\leq L\sqrt{2}\mathbb{E}_{\sigma^{\prime}}\left[\sup_{x\in\mathrm{CH}(\mathcal{X})}\sum_{i=1}^{n}x_{i}\sum_{k=1}^{K}\sigma^{\prime}_{ik}\right],$
where $(\sigma_{k})_{k\in[K]}$ and $(\sigma^{\prime}_{ki})_{(k,i)\in[K]\times[n]}$ are independent Rademacher variables.
Proposition 4.11 (Corollary D.11 from Mohri et al. (2018))
$\mathbb{E}_{\sigma^{\prime}}\left[\max_{i\in[n]}\sum_{k=1}^{K}\sigma^{\prime}_{ik}\right]\leq\sqrt{2K\log n},$
where $(\sigma^{\prime}_{ki})_{(k,i)\in[K]\times[n]}$ are independent Rademacher variables.
We now show that our objective function of interest is indeed Lipschitz in the placement $x$.
Lemma 4.12
Functions $(h_{k}(x))_{k\in[K]}$ defined as $h_{k}(x)=1-\mathsf{OFF}(x,D^{k})/Q$ are $(\sqrt{n}/Q)$-Lipschitz in the $\ell_{2}$ norm.
The proof of this lemma is deferred to Subsection 6.7. With this we are finally able to write the proof of Theorem 4.8, which is presented in Subsection 6.8.
4.4 Approximation Algorithm and Main Result
We now have all the tools required to prove the guarantees for our coordinated placement and fulfillment procedures. To summarize the procedures, we first sample $K$ independent demand realizations $D^{1},\dots,D^{K}$. Then, we solve for fractional placement $\hat{x}$, an optimal solution of $\max_{x\in\mathrm{CH}(\mathcal{X})}\widehat{\mathsf{OFF}}(x)$. Once we have $\hat{x}$, we will feed it to our randomized rounding procedure and obtain an integer placement $R(\hat{x})$. We then use the fulfillment procedure that is appropriate to the demand model (either Spatial or Temporal Independence) to find a policy $\pi$ that is approximately-optimal given the placement $R(\hat{x})$. Overall, we obtain the following guarantee.
Theorem 4.13
For any instance where the maximum number of DCs that can serve a single demand type is $d$, we can efficiently obtain an inventory placement $R(\hat{x})$ and a fulfillment policy $\pi$ with
$\mathbb{E}_{\wedge,R}\left[\mathsf{ONL}^{\pi}(R(\hat{x}))\right]\geq\frac{1}{2}\left(1-\left(1-\frac{1}{d}\right)^{d}\right)\max_{x\in\mathcal{X}}\mathsf{OFF}(x)-\varepsilon,$
using $K=O\left(\frac{Q^{2}n\log n}{\varepsilon^{2}}\right)$ demand samples.
The proof of this theorem is deferred to Subsection 6.9.
We finish this section by highlighting the practical aspects of our placement procedure in Theorem 4.13. First, it is completely sample-driven. In practice, past demand can be used as demand samples, and thus there is no need to fit an exact demand distribution. Second, our procedure only required samples of aggregate demand over the time horizon, so it is not reliant on knowing the order in which each customer from each location arrived. Third, it is straightforward to obtain by solving a linear program and potentially needing to round the solution. We say potentially because we have found that having a fractional solution $\hat{x}$ seems to be a rare occurrence. Indeed, $\hat{x}$ turned out to be integral in all of our experiments based on real-world data. In all cases in which $\hat{x}$ is integral, our guarantee improves to $1/2$, with the only loss coming from fulfillment and sampling.
All in all, Theorem 4.13 formalizes the notion that assuming offline fulfillment has a good outcome if a good fulfillment policy is being deployed. To elaborate, our placement procedure, which solves the offline relaxation to determine the inventory placement, is optimizing as if the downstream fulfillment team can make decisions with perfect offline hindsight. Our result shows that this leads to state-of-the-art approximation guarantees when the fulfillment procedure is indeed quite good, in the sense of having a competitive ratio relative to offline. This intuition makes sense—a good fulfillment team means that the placement team can be optimistic when making assumptions on how well the placed inventory will be distributed.
5 Experimental Results
Generally speaking, an instance of the e-commerce placement and fulfillment problem consists of a quantity $Q$ to be placed across DCs, a network defined by rewards $r_{ij}$ between DCs $i$ and demand locations $j$, and a distribution over arrival sequences from different demand locations. In Section 4 we analyzed worst-case instances under theoretical structural forms (Spatial and Temporal Independence) for the distribution over arrival sequences. In this section we extract realistic distributions over arrival sequences from publicly-available JD.com data (Shen et al. 2020), and also consider simplified reward topologies motivated by the JD.com network.
5.1 Description of JD.com Network
JD.com is a large Chinese e-tailer that separates its territory into geographical regions. Each geographical region is further divided into districts, and each district comes with a distribution center. A singular district in each region has a Regional Distribution Center (RDC): a big DC that usually holds more inventory and can fulfill demand to any district within the region. The other districts within a region are equipped with a Front Distribution Center (FDC): a smaller DC primarily dedicated to fulfilling demand from within the district.
For our experiments, we will assume that FDCs can only fulfill demand locally. The RDC will be able to fulfill the demand from any district in the region. Formally, we use $i\in[n]_{0}:=\{0,\dots,n\}$ to denote districts. District 0 corresponds to the one equipped with the RDC, and can thus fulfill demand to any other district. Since we don’t have information about fulfillment costs or geographical coordinates, we will assume that demand fulfilled locally yields a reward of 1. (The reward of fulfilling locally at district $i=0$ is slightly inflated to $1+10^{-7}$ so that inventory placement favors local fulfillment at the RDC over FDC’s.). Any spillover demand (which can only occur from district 0 to some district $i\neq 0$) gives a reward of $r\in(0,1)$.
This network structure simplifies the fulfillment problem to an accept/reject problem. (In practice, a rejection can be interpreted as delaying a fulfillment long enough so that there is inventory in the corresponding FDC to fulfill it.) Whenever a demand request that can be fulfilled locally arrives, any policy should agree to fulfill it. The challenging case is when a demand request from district $i\neq 0$ arrives, there is no inventory left at FDC $i$, and there is remaining inventory at the RDC. The e-tailer is faced with the decision of either accepting the spillover fulfillment and collecting a reward of $r$, or rejecting it to save the unit of inventory and potentially using it for a local fulfillment at 0, collecting a reward of $1>r$. Inspired by the statistical hypothesis testing nomenclature, we say that the e-tailer makes a type 1 error when they reject a spillover order that could have been fulfilled without hurting local fulfillments at District 0. On the other hand, we say that the e-tailer makes a type 2 error when they fulfill a spillover order using a unit of inventory that could have been saved for a local fulfillment.
5.2 Data Description and Pre-processing
The dataset contains information about 549,989 completed orders of 31,868 different SKUs during March of 2018. An order can be associated with multiple SKUs. For each SKU in an order, we have information about the moment it was checked out, to which district the order must be delivered, from which district it was delivered, and the number of units that were ordered.
The information extracted from the dataset is about demand patterns. The dataset does contain information about paid prices, but since we do not know any cost information, we assume that reward depends only on whether fulfillment was done locally or through a spillover.
For our experiments, we focus on the three regions with the highest total demand: regions 9, 2, and 5. We divide our demand data by week, and the logic behind this is to think about weekly inventory replenishment. Since the data consists of a month, we only have three full weeks, and we discard any order not made during these 3 weeks. We also separate the orders by SKU, ignoring how orders containing multiple items are split. For each SKU-week pair, we take the stream of demand requests from the data. Each request indicates the district from which it is originating and a time stamp. If an order contains more than 1 unit, we create duplicate demand requests that arrive at the same time. The first two full weeks of the month are considered the training set and will be used to determine inventory placements and fulfillment policies. The third week is the test set and will serve to evaluate the different procedures for determining inventory placements and fulfillment policies.
Since we only have 3 weeks of data, we pool SKUs with similar demand behavior to enlarge both our training and testing data. For each region, we select SKUs that have an average weekly demand between 20 and 40 units across the three weeks. We further impose that the coefficient of variation of the weekly demand is not larger than 0.5. With this way of filtering SKUs we seek to accomplish two goals. The first one is to obtain weeks with similar, stationary demand patterns so that the demand behavior of the training set serves to predict the demand patterns to appear in the test set. The second one is to filter out censored demand numbers, since these will generally have a large coefficient of variation where the observed sales are high and then suddenly drops to 0 once inventory runs out.
5.3 Placement Procedures
In our experiments, a placement procedure is given $Q$ and all values of $r_{ij}$, and must decide how to split $Q$ units of inventory among DCs. Importantly, a placement procedure does not know the true distribution over arrival sequences on which it will be tested; instead, it is given as training samples the arrival sequences for the same items from chronologically-earlier weeks, as described in Subsection 5.2. Let $k=1,\ldots,K$ denote the training arrival sequences. For all $i\in[n]_{0}$ and $k\in[K]$, we let $D^{i}_{k}$ denote the total demand from district $i$ in sample $k$. We let $\bar{D}_{i}=\sum_{k=1}^{K}D_{i}^{k}/K$ denote the average demand from district $i$ over the samples.
We consider four placement procedures: Proportional, Offline, Fluid, and Myopic. Proportional and Fluid placement make use of only the values of $\bar{D}_{i}$. Offline placement, which was analyzed in Section 4, makes use of the entire distribution of aggregate demands $(D^{k}_{i})_{k\in[K]}$ from each district $i$. Myopic runs the exact arrival sequences through a simulator to determine the best placement. We note that all of these are standard placement procedures (see DeValve et al. 2023b).
5.3.1 Proportional Placement.
This is the simplest placement procedure. It simply places the inventory proportional to the expected demand in each district. This placement sets $x_{i}=Q\bar{D}_{i}/\sum_{j=0}^{n}\bar{D}_{j}$, which is then rounded arbitrarily to obtain $x^{\mathsf{PROP}}$ while satisfying $\sum_{i\in[n]_{0}}x^{\mathsf{PROP}}_{i}=Q$.
5.3.2 Offline Placement.
This placement procedure assumes that the fulfillment will be done in a hindsight optimal or offline manner. It solves the following LP, with variables $(x_{i})_{i\in[n]_{0}}$ for deciding the inventory in district $i$, variables $(y_{i}^{k})_{i\in[n]_{0},k\in[K]}$ to decide the local fulfillment at district $i$ in sample $k$, and variables $(z_{i}^{k})_{i\in[n],k\in[K]}$ for deciding the spillover fulfillment from the RDC to district $i\neq 0$ in sample $k$. The LP is the following:
$(\mathsf{OFF})\quad\max_{x,y,z}\quad$ $\frac{1}{K}\sum_{k=1}^{K}\left[\sum_{i\in[n]_{0}}y_{i}^{k}+r\sum_{i\in[n]}z_{i}^{k}\right]$
$\mathrm{s.t.}\quad$ $\sum_{i\in[n]_{0}}x_{i}=Q$
$y_{i}^{k}+z_{i}^{k}\leq D_{i}^{k}$ $\forall i\in[n],k\in[K]$
$y_{0}^{k}\leq D_{0}^{k}$ $\forall k\in[K]$
$y_{i}^{k}\leq x_{i}$ $\forall i\in[n],k\in[K]$
$y_{0}^{k}+\sum_{i\in[n]}z_{i}^{k}\leq x_{0}$ $\forall k\in[K]$
$x,y,z\geq 0.$
We note that $\mathsf{OFF}$ is identical to the sample-average approximation LP from Subsection 4.3 if the network is specialized to the JD.com network. After the LP is solved, the optimal solution $x$ is randomly rounded to obtain $x^{\mathsf{OFF}}$ while satisfying $\sum_{i\in[n]_{0}}x^{\mathsf{OFF}}_{i}=Q$. Remarkably, all of the solutions to the above linear program turned out to be integers, so no rounding was needed.
5.3.3 Fluid Placement.
This placement procedure is similar to the Offline placement, but it solves a fluid approximation of the stochastic linear program. The variables $(x_{i})_{i\in[n]_{0}}$, $(y_{i})_{i\in[n]_{0}}$ and $(z_{i})_{i\in[n]}$ have the same interpretation, but they are no longer indexed by the sample $k$. The LP is as follows:
$(\mathsf{FLU})\quad\max_{x,y,z}\quad$ $\sum_{i\in[n]_{0}}y_{i}+r\sum_{i\in[n]}z_{i}$
$\mathrm{s.t.}\quad$ $\sum_{i\in[n]_{0}}x_{i}=Q$
$y_{i}+z_{i}\leq\bar{D}_{i}$ $\forall i\in[n]$
$y_{0}\leq\bar{D}_{0}$
$y_{i}\leq x_{i}$ $\forall i\in[n]$
$y_{0}+\sum_{i\in[n]}z_{i}\leq x_{0}$
$x,y,z\geq 0.$
After the LP is solved, the optimal solution $x$ is rounded arbitrarily to obtain $x^{\mathsf{FLU}}$ while satisfying $\sum_{i\in[n]_{0}}x^{\mathsf{FLU}}_{i}=Q$.
5.3.4 Myopic Placement.
This placement procedure assumes that a greedy fulfillment policy will be deployed. In this context, greedy means that if there is inventory available to fulfill locally, demand is fulfilled locally. If there is no inventory to fulfill locally, but there is inventory available at the RDC, the demand will be satisfied through spillover fulfillment. Demand originating from district $i$ is only lost if there is no leftover inventory at the FDC and RDC (only at the RDC if $i=0$). The placement $x^{\mathsf{MYO}}$ is obtained through a simulation optimization approach, performing iterative local improvements which consist of moving a single unit of inventory from one DC to another. This pair of DCs is chosen so that the marginal gain in expected reward after each move is maximized.
5.4 Fulfillment Procedures
A fulfillment procedure is given an inventory placement and all values of $r_{ij}$, and must determine a fulfillment policy that dynamically decides on a DC to serve each arriving demand. Like before, the fulfillment procedure is only given training samples that are representative of the test sequence.
We consider three fulfillment procedures: Myopic, Fluid Shadow Prices, and Stochastic Shadow Prices, with there being two versions of each shadow price policy—the ‘static’ one where shadow prices are fixed from the beginning, and a ‘re-solving’ one where shadow prices are updated daily. Again, all of these are standard fulfillment procedures (see DeValve et al. 2023b). We do not computationally test the fulfillment policies used to establish Theorem 4.13, since they are designed for the Spatial and Temporal Independence stochastic models which are not imposed in our experiments.
5.4.1 Myopic Fulfillment.
This procedure always outputs a myopic fulfillment policy. As mentioned earlier, this policy corresponds to using a greedy algorithm to make fulfillment decisions. Refer to 5.3.4 for a description.
5.4.2 Fluid Shadow Prices (F-SP).
This procedure uses the shadow prices of the linear program $(\mathsf{FLU})$ to account for the opportunity cost of performing a spillover fulfillment. Specifically, let $\lambda$ be an optimal dual variable for the constraint
$y_{0}+\sum_{i\in[n]}z_{i}\leq x_{0}.$
This dual variable is used to approximate the opportunity cost of depleting a unit of inventory from the RDC. Thus, the resulting policy is as follows. If there is enough inventory to fulfill locally, do so. If there is not enough inventory to fulfill locally and there is inventory at the RDC, perform a spillover fulfillment if and only if $r>\lambda$.
5.4.3 Stochastic Shadow Prices (S-SP).
This procedure follows the same logic as the previous one but uses the linear program $(\mathsf{OFF})$ to account for the opportunity costs instead. For each $k\in[K]$ let $\lambda_{k}$ be the optimal dual variable associated with the constraint
$y_{0}^{k}+\sum_{i\in[n]}z_{i}^{k}\leq x_{0},$
and let $\lambda=\sum_{k=1}^{K}\lambda_{k}$. The description of the policy is then identical to that for Fluid Shadow Prices.
5.5 Experimental Setup
We have 3 regions, each defining a (training and test) distribution over arrival sequences. For any region we can attach values of $Q$ and $r$ to create an instance. On each instance, we can run one of the 4 placement procedures from Subsection 5.3, and based on each placement, run one of the 5 fulfillment procedures from Subsection 5.4. We also add a 6th fulfillment procedure that always makes the hindsight-optimal offline fulfillments, as a hypothetical benchmark.
To make performances comparable across instances, we always divide by the reward of an omniscient who can perform fulfillment offline, and moreover knows the true test (not training) distribution and is allowed to make a fractional inventory placement. Formally, the reward of this omniscient is defined by the quantity $\max_{x\in\mathrm{CH}(\mathcal{X})}\mathsf{OFF}(x)$ from Section 4, fed with the empirical distribution corresponding to the test set. For a combination of placement and fulfillment procedure, we refer to the ratio of its reward to that of this omniscient as the competitive ratio, on a given instance.
For each region, we create instances with $r$ ranging in $R:=\{0.1,0.5,0.9\}$, and $Q$ ranging orthogonally so that the load factor is as close as possible to one of the 9 values in $\{0.5,0.75,1,1.25,1.5,1.75,2,2.25,2.5\}$. The load factor is defined as the quotient between the expected weekly demand of a region and the total inventory to be placed (this expected demand is computed considering both the train and test sets).
5.6 Results
We have 24 combinations of placement and fulfillment procedures to test (4 placement procedures, multiplied by 6 fulfillment procedures including offline fulfillment). Each combination is tested on 81 instances (3 regions, multiplied by 3 values for $r$, multiplied by 9 values for $Q$). We first report the average competitive ratios for each combination over the 27 instances with the same value of $r$, in Table 2. Across the columns, OFF, MYO, PROP, and FLU stand for Offline, Myopic, Proportional, and Fluid Placement, respectively; down the rows, Myopic, F-SP-s, F-SP-r, S-SP-s, S-SP-r, and Offline stand for Myopic, Fluid Shadow Prices that are static or re-solving, Stochastic Shadow Prices that are static or re-solving, and Offline Fulfillment, respectively. We note that OFF placement combined with Offline fulfillment does not achieve 100%, because the omniscient has the further benefits of knowing the test distribution and being allowed to make a fractional placement.
$r=0.1$ $r=0.5$ $r=0.9$
OFF MYO PROP FLU OFF MYO PROP FLU OFF MYO PROP FLU
Myopic 88.95 92.04 91.18 76.77 92.59 94.09 93.65 87.64 98.17 98.24 94.38 96.37
F-SP-s 98.63 97.11 97.44 92.42 96.24 96.55 95.85 91.77 88.62 89.29 92.77 89.65
F-SP-r 98.33 96.98 97.34 91.74 97.40 97.24 96.31 93.78 95.69 95.81 93.73 94.15
S-SP-s 98.89 97.36 97.69 92.58 96.44 96.69 95.85 91.78 96.96 97.35 94.02 95.35
S-SP-r 98.78 97.33 97.66 92.17 97.59 97.39 96.40 93.91 98.57 98.67 94.68 96.84
Offline 99.49 97.72 98.10 93.77 99.71 98.73 97.40 96.86 99.87 99.83 95.10 98.14
Table 2: Competitive ratio for different placement-fulfillment pairs, averaged over all load factors and regions. For each value of $r$ and fulfillment procedure, the highest competitive ratio among placement procedures is bolded.
When comparing the performance of different placement procedures, the first thing we notice is that, as expected, Myopic Placement is the best alternative if Myopic Fulfillment is being deployed, and Offline Placement has better performance if Offline Fulfillment is being deployed. For the remaining placement procedures, we break down the performance depending on $r$. If $r=0.1$, then Offline Placement consistently dominates the performance of the other placement procedures. When $r=0.5$, Offline Placement and Myopic Placement have very similar performance, which is again better than the performance obtained by the other two placement procedures. When $r=0.9$, then Myopic Placement tends to be slightly better than Offline Placement. An explanation for the superiority of Offline and Myopic Placement is that they take the stochastic nature of demand into account by maximizing the average reward in various demand scenarios. On the other hand, Fluid and Proportional placement only use information about the average demand. When comparing Myopic and Offline placements, the latter has an ‘informational disadvantage’ against the former, as Offline placement does not use information about the demand arrival order. Even so, Offline Placement performs as well or better than Myopic Placement. Offline Placement also has a computational advantage over Myopic Placement, since the former can be solved through a linear program (up to rounding errors). Myopic Placement, on the other hand, has to be obtained through simulation-based optimization and local search, which is costly and implementation-sensitive.
We also observe that S-SP with re-solving achieves the highest or second-highest average competitive ratio among all fulfillment procedures, for any placement procedure, and when it is second highest, it is only by a small margin. Moreover, the fulfillment procedures whose performance is close to the one obtained by S-SP with re-solving vary with $r$. For $r=0.1$, the static shadow price policies (F-SP-s and S-SP-s) perform well. This makes sense since not re-solving will induce a higher threshold for performing spillover fulfillment, which yields a low reward. In other words, type 1 errors are not very costly, so good algorithms are the ones that prioritize not making type 2 errors. For $r=0.5$ the policy closest to S-SP-r is F-SP-r. In this regime, spillover rewards give exactly half the reward of local rewards, so good algorithms are the ones that carefully balance the two types of errors. These algorithms are the ones that re-solve. Finally, for $r=0.9$, the fulfillment procedure that performs closely to S-SP with re-solving is Myopic Fulfillment. This is the mirror case of $r=0.1$, so good fulfillment policies are the ones that prioritize not making type 1 errors. This is precisely Myopic Fulfillment, which never makes type 1 errors (but is very prone to make type 2 errors). All in all, the results suggest that S-SP with re-solving is a robust fulfillment algorithm that achieves the best performance irrespective of the placement procedure and value of $r$.
Figure 1: Performance of S-SP with re-solving with different inventory placements for $r\in\{0.1,0.5,0.9\}$.
Since S-SP with re-solving dominates, we hereafter fix it to be the fulfillment procedure, which allows us to further distinguish performance along other dimensions. In Figure 1 we plot the competitive ratio for different placement procedures and different values of $r$ further distinguished by load factor, averaged over the three studied regions. We observe that for $r\in\{0.5,0.9\}$, it is hard to distinguish which one is better between Offline Placement or Myopic Placement. This also happens when $r=0.1$ with load factors lower or equal to 1. If $r=0.1$ and the load factor is at least 1.25, then Offline Placement exhibits consistently better performance. Based on these results, we argue that Offline Placement seems to be the most robust placement procedure out of the four when a high-quality fulfillment policy is being deployed.
One last thing to analyze is the performance of Fluid Placement. For $r\in\{0.1,0.5\}$ we observe the same pattern—that performance peaks at load factors near 1, and decreases as it gets away from that value. This likely happens because Fluid Placement is prioritizing placement at the RDC, regardless of the value of $r$. (The non-dependence on the value of $r$ is explained by the deterministic nature of the Fluid LP which allows to construct optimal solutions using a greedy approach.) For load factors higher than 1 (more expected demand than inventory) the preference of the RDC over the FDCs can be explained because of the $10^{-7}$ perturbation added to the local reward at District 0. (Fluid Placement only performs worse if we tie-break in favor of FDC placement. This is especially true in settings where the value of $r$ is large.) This will make an optimal solution for the fluid LP to satisfy all of the demand from District 0 before it starts placing inventory at the FDCs. There is an improvement in performance when the load factor exceeds 1.75 (very scarce inventory), where placing inventory solely at the RDC is a good solution since all of the scarce units will likely be sold. For load factors lower than 1 (more inventory than expected demand) the problem becomes degenerate in the sense that it has multiple solutions, as all demand is saturated once $Q$ equals the expected demand and leftover inventory won’t contribute to the objective. We also tiebreak in favor of the RDC, which explains the comparably good performance when $r=0.9$. We conclude that overall, Fluid Placement is unstable and suffers from multiple optimal solutions, and performs worse than the other placement procedures irrespective of tiebreaking.